a total of $8000 is invested into two simple interest accounts. On one account the annual simple interest rate is 9%; on the second account the annual simple interest rate is 7% How much should be invested in the 7% account so that the total interest earned is 8.5% of the total investment?
Answer by mananth(16949)

a total of $8000 is invested into two simple interest accounts. On one account the annual simple interest rate is 9%; on the second account the annual simple interest rate is 7% How much should be invested in the 7% account so that the total interest earned is 8.5% of the total investment?
.
Let $x be the investment in 7% account
8000-x will be invested in 9%
Total invested = $ 8000
you should get 8.5%
..
0.07x +0.09(8000-x)=8000*0.085
0.07x+720-0.09x=680
-0.02x=680-720
x=-40/-0.02
x= $2000